# label*label*footnotetext: Corresponding authors

### v1

@@ Abstract
This note aims to present novel positive linear operators involving the Wright function. Furthermore, the present research established the moments of these newly defined operators and estimated the convergence rate using the classical modulus of continuity. Additionally, the convergence rate in the Lipschitz spaces, their $A$-statistical convergence property, and $(\lambda,\gamma)-$ statistical convergence properties have been covered.
@@
2000 Mathematics Subject Classification: primary 41A25, 41A30, 41A36.
journal: arXiv
@@ 1 Introduction
A new generalization of Bernstein operators was introduced periodically in approximation theory (see totik1994approximation ; lewanowicz2004generalized ; sinha2009inverse ; gupta2013approximation ; gupta2021operators ; mishra2014generalized ). Many positive linear operators are derived from special functions. For instance, the operators were introduced and studied in ozarslan2013astatistical using the Mittag-Leffler function, the operators were studied in gupta2020convergence with the help of gamma function, and using Laguerre polynomials, the positive linear operators introduced by Gupta gupta2024new . This note aims to merge two mathematical topics-approximation theory and special functions.
@@
named after the British mathematician E. M. Wright, has appeared for the first time in the case $\rho>0$ in connection with his investigations in the asymptotic theory of partitions in wright1933coefficients .
Note that $\phi(1,1;z)=I_{0}(2\sqrt{z})$, where $I_{n}(z)=i^{-n}J_{n}(iz)=\dfrac{(z/2)^{n}}{\Gamma(1+n)}{}_{0}F_{1}\left(-;1+n;\dfrac{z^{2}}{4}\right)$, $n$ not a negative integer. The function $I_{n}(z)$ is called the modified Bessel function of the first kind of index $n$. Here, $J_{n}$ is Bessel function and ${}_{0}F_{1}$ is hypergeometric function. The modified function $I_{n}$ is related to $J_{n}$ in much the same way that the hyperbolic function is related to the trigonometric function.
@@
It was demonstrated in gray1895treatise ; bateman1953higher that, the modified Bessel function $I_{n}(z)$ of the first kind does not have zeros in the right half-plane. khanmamedov2010zeros contains an analysis of $n$-zeros of the function $I_{n}(z)$ in the left half-plane. Since $I_{n}(z)$ has no zeros in $[0,\infty)$, the Wright function $\phi_{1,m}(z)$ also has no zeros in $[0,\infty)$.
@@
$W_{n}^{(\beta)}(f;x)=\dfrac{1}{\phi_{1,\beta}\left(nx\right)}\sum_{k=0}^{\infty}f\left(\dfrac{k}{n}\right)\dfrac{(nx)^{k}}{k!\Gamma(k+\beta)},$ (1.2)
where $f\in E:=\left\{f\in C\left[0,\infty\right):\lim_{x\to\infty}\dfrac{f(x)}{1+x^{2}}\textrm{ is finite }\right\}$ and $C\left[0,\infty\right)$ denote the space of continuous functions defined on $[0,\infty)$. Recall that the Banach lattice $E$ is endowed with the norm
@@
Research on approximating continuous signals using a sequence of positive linear operators is ongoing. We provide other findings along the same lines in this note, but with the aid of recently established positive linear operators $W_{n}^{(\beta)}$. There is no prior literature that introduced these operators (1.2). We estimate the moments and central moments of the operators (1.2) up to the fourth order in the second section. The rate of convergence of these positive linear operators is discussed in the next section. In the same section 3, we have proved that $W_{n}^{(\beta)}$ maps $E$ into itself. We have also determined the rate of convergence for these operators by utilizing the modulus of continuity. For the operators $W_{n}^{(\beta)}$, we have established a statistical Voronovskaya-type theorem in section 4. A weighted $\lambda\beta$-statistical approximation theorem for the Wright operators (1.2) has been established in section 5. It should be noted that depending on the choice of $\lambda$ and $\gamma$, $\lambda\gamma$-statistical convergence is a non-trivial extension of ordinary and statistical convergence. Some extensions of the Wright operators are shown in the last part to estimate the functions in $L_{p}$ space for the next research.
@@ 2 Some Lemmas
Lemma 1
(mehrez2017functional, , Theorem 6.1) Let $\alpha,\beta>0$. Then the following assertions are true:
@@
    $\left|W_{n}^{(\beta)}(t;x)-x\right|\leq 0;$
@@
    $\left|W_{n}^{(\beta)}(t^{2};x)-x^{2}\right|\leq\dfrac{x}{n\beta};$
@@
    $\left|W_{n}^{(\beta)}(t^{3};x)-x^{3}\right|\leq\dfrac{3x^{2}}{n\beta(\beta+1)}+\dfrac{x}{n^{2}\beta};$
@@
    $\left|W_{n}^{(\beta)}(t^{4};x)-x^{4}\right|\leq\dfrac{6x^{3}}{n\beta(\beta+1)(\beta+2)}+\dfrac{7x^{2}}{n^{2}\beta(\beta+1)}+\dfrac{x}{n^{3}\beta}.$
Proof: Since,
@@
$W_{n}^{(\beta)}(t;x)$ $=\dfrac{1}{\phi_{1,\beta}\left(nx\right)}\sum_{k=1}^{\infty}\dfrac{1}{n}\dfrac{(nx)^{k}}{(k-1)!\Gamma(k+\beta)}$
$=\dfrac{x}{\phi_{1,\beta}\left(nx\right)}\sum_{k=0}^{\infty}\dfrac{(nx)^{k}}{k!\Gamma(k+\beta+1)}=x\dfrac{\phi_{1,\beta+1}(nx)}{\phi_{1,\beta}(nx)}\leq\dfrac{x}{\beta}\leq x\textrm{ for }\beta>1.$
@@
$W_{n}^{(\beta)}(t^{2};x)$ $=\dfrac{1}{\phi_{1,\beta}\left(nx\right)}\sum_{k=1}^{\infty}\dfrac{(k-1)+1}{n^{2}}\dfrac{(nx)^{k}}{(k-1)!\Gamma(k+\beta)}$
$=\dfrac{x^{2}}{\phi_{1,\beta}\left(nx\right)}\sum_{k=0}^{\infty}\dfrac{(nx)^{k}}{k!\Gamma(k+\beta+2)}+\dfrac{x}{n\phi_{1,\beta}\left(nx\right)}\sum_{k=0}^{\infty}\dfrac{(nx)^{k}}{k!\Gamma(k+\beta+1)}$
$=x^{2}\dfrac{\phi_{1,\beta+2}\left(nx\right)}{\phi_{1,\beta}\left(nx\right)}+\dfrac{x}{n}\dfrac{\phi_{1,\beta+1}\left(nx\right)}{\phi_{1,\beta}\left(nx\right)}\leq\left|\dfrac{x^{2}\Gamma(\beta)}{\Gamma(\beta+2)}\right|+\left|\dfrac{x\Gamma(\beta)}{n\Gamma(\beta+1)}\right|=\dfrac{x^{2}}{\beta(\beta+1)}+\dfrac{x}{\beta n}.$
@@
$\left|W_{n}^{(\beta)}(t^{2};x)-x^{2}\right|\leq\dfrac{x}{\beta n}.$
@@
$W_{n}^{(\beta)}(t^{3};x)$ $=$ $\dfrac{1}{\phi_{1,\beta}\left(nx\right)}\sum_{k=1}^{\infty}\dfrac{k^{2}}{n^{3}}\dfrac{(nx)^{k}}{(k-1)!\Gamma(k+\beta)}$
$=$ $\dfrac{1}{\phi_{1,\beta}\left(nx\right)}\left[\sum_{k=2}^{\infty}\dfrac{k+1}{n^{3}}\dfrac{(nx)^{k}}{(k-2)!\Gamma(k+\beta)}+\sum_{k=1}^{\infty}\dfrac{1}{n^{3}}\dfrac{(nx)^{k}}{(k-1)!\Gamma(k+\beta)}\right]$
$=$ $\dfrac{1}{\phi_{1,\beta}\left(nx\right)}\left[\sum_{k=3}^{\infty}\dfrac{1}{n^{3}}\dfrac{(nx)^{k}}{(k-3)!\Gamma(k+\beta)}+\sum_{k=2}^{\infty}\dfrac{3}{n^{3}}\dfrac{(nx)^{k}}{(k-2)!\Gamma(k+\beta)}+\sum_{k=1}^{\infty}\dfrac{1}{n^{3}}\dfrac{(nx)^{k}}{(k-1)!\Gamma(k+\beta)}\right]$
$=$ $x^{3}\dfrac{\phi_{1,\beta+3}\left(nx\right)}{\phi_{1,\beta}\left(nx\right)}+3x^{2}\dfrac{\phi_{1,\beta+2}\left(nx\right)}{n\phi_{1,\beta}\left(nx\right)}+x\dfrac{\phi_{1,\beta+1}\left(nx\right)}{n^{2}\phi_{1,\beta}\left(nx\right)}$
$\leq$ $\left|\dfrac{x^{3}\Gamma(\beta)}{\Gamma(\beta+3)}\right|+\left|\dfrac{3x^{2}\Gamma(\beta)}{n\Gamma(\beta+2)}\right|+\left|\dfrac{x\Gamma(\beta)}{n^{2}\Gamma(\beta+1)}\right|=\dfrac{x^{3}}{\beta(\beta+1)(\beta+2)}+\dfrac{3x^{2}}{n\beta(\beta+1)}+\dfrac{x}{\beta n^{2}}.$
@@
$\left|W_{n}^{(\beta)}(t^{3};x)-x^{3}\right|\leq\dfrac{3x^{2}}{n\beta(\beta+1)}+\dfrac{x}{\beta n^{2}}.$
@@
$W_{n}^{(\beta)}(t^{4};x)$ $=$ $\dfrac{1}{\phi_{1,\beta}\left(nx\right)}\sum_{k=1}^{\infty}\dfrac{k^{3}-1+1}{n^{4}}\dfrac{(nx)^{k}}{(k-1)!\Gamma(k+\beta)}$
$=$ $\dfrac{1}{\phi_{1,\beta}\left(nx\right)}\left[\sum_{k=2}^{\infty}\dfrac{k^{2}+k+1}{n^{4}}\dfrac{(nx)^{k}}{(k-2)!\Gamma(k+\beta)}+\sum_{k=1}^{\infty}\dfrac{1}{n^{4}}\dfrac{(nx)^{k}}{(k-1)!\Gamma(k+\beta)}\right]$
$=$ $\dfrac{1}{\phi_{1,\beta}\left(nx\right)}\left[\sum_{k=2}^{\infty}\dfrac{k^{2}-4+4}{n^{4}}\dfrac{(nx)^{k}}{(k-2)!\Gamma(k+\beta)}+\sum_{k=2}^{\infty}\dfrac{k+1}{n^{4}}\dfrac{(nx)^{k}}{(k-2)!\Gamma(k+\beta)}+\sum_{k=1}^{\infty}\dfrac{1}{n^{4}}\dfrac{(nx)^{k}}{(k-1)!\Gamma(k+\beta)}\right]$
$=$ $\dfrac{1}{n^{4}\phi_{1,\beta}\left(nx\right)}\left[\sum_{k=3}^{\infty}\dfrac{(k+2)(nx)^{k}}{(k-3)!\Gamma(k+\beta)}+\sum_{k=2}^{\infty}\dfrac{4(nx)^{k}}{(k-2)!\Gamma(k+\beta)}+\sum_{k=2}^{\infty}\dfrac{(k-2+3)(nx)^{k}}{(k-2)!\Gamma(k+\beta)}+\sum_{k=1}^{\infty}\dfrac{(nx)^{k}}{(k-1)!\Gamma(k+\beta)}\right]$
$=$ $\dfrac{1}{n^{4}\phi_{1,\beta}\left(nx\right)}\left[\sum_{k=4}^{\infty}\dfrac{(nx)^{k}}{(k-4)!\Gamma(k+\beta)}+\sum_{k=2}^{\infty}\dfrac{7(nx)^{k}}{(k-2)!\Gamma(k+\beta)}+\sum_{k=3}^{\infty}\dfrac{6(nx)^{k}}{(k-3)!\Gamma(k+\beta)}+\sum_{k=1}^{\infty}\dfrac{(nx)^{k}}{(k-1)!\Gamma(k+\beta)}\right]$
$=$ $x^{4}\dfrac{\phi_{1,\beta+4}\left(nx\right)}{\phi_{1,\beta}\left(nx\right)}+6x^{3}\dfrac{\phi_{1,\beta+3}\left(nx\right)}{n\phi_{1,\beta}\left(nx\right)}+7x^{2}\dfrac{\phi_{1,\beta+2}\left(nx\right)}{n^{2}\phi_{1,\beta}\left(nx\right)}+x\dfrac{\phi_{1,\beta+1}\left(nx\right)}{n^{3}\phi_{1,\beta}\left(nx\right)}$
$\leq$ $\left|\dfrac{x^{4}\Gamma(\beta)}{\Gamma(\beta+4)}\right|+\left|\dfrac{6x^{3}\Gamma(\beta)}{n\Gamma(\beta+3)}\right|+\left|\dfrac{7x^{2}\Gamma(\beta)}{n^{2}\Gamma(\beta+2)}\right|+\left|\dfrac{x\Gamma(\beta)}{n^{3}\Gamma(\beta+1)}\right|$
$=$ $\dfrac{x^{4}}{\beta(\beta+1)(\beta+2)(\beta+3)}+\dfrac{6x^{3}}{n\beta(\beta+1)(\beta+2)}+\dfrac{7x^{2}}{n^{2}\beta(\beta+1)}+\dfrac{x}{n^{3}\beta}.$
@@
$\left|W_{n}^{(\beta)}(t^{4};x)-x^{4}\right|\leq\dfrac{6x^{3}}{n\beta(\beta+1)(\beta+2)}+\dfrac{7x^{2}}{n^{2}\beta(\beta+1)}+\dfrac{x}{n^{3}\beta}.$
Lemma 3
@@
    $W_{n}^{(\beta)}(\phi_{x}^{2};x)\leq\dfrac{x}{n\beta}$
@@
    $W_{n}^{(\beta)}(\phi_{x}^{3};x)\leq\dfrac{3x^{2}(\beta+2)}{n\beta(\beta+1)}+\dfrac{x}{n^{2}\beta}.$
@@
    $W_{n}^{(\beta)}(\phi_{x}^{4};x)\leq\dfrac{6x^{3}\left(\beta^{2}+5\beta+7\right)}{n\beta(\beta+1)(\beta+2)}+\dfrac{4x^{2}\beta+11}{n^{2}\beta(\beta+1)}+\dfrac{x}{n^{3}\beta}.$
Proof: Using Lemma 2 and simple computation gives $W_{n}^{(\beta)}(\phi_{x}^{1};x)=0$ and
$W_{n}^{(\beta)}(\phi_{x}^{2};x)\leq\left|W_{n}^{(\beta)}(t^{2};x)-x^{2}\right|+2x\left|W_{n}^{(\beta)}(t;x)-x\right|+x^{2}\left|W_{n}^{(\beta)}(1;x)-1\right|\leq\dfrac{x}{n\beta}.$
@@
$\leq\dfrac{6x^{3}}{n\beta(\beta+1)(\beta+2)}+\dfrac{7x^{2}}{n^{2}\beta(\beta+1)}+\dfrac{x}{n^{3}\beta}+4x\left(\dfrac{3x^{2}}{n\beta(\beta+1)}+\dfrac{x}{n^{2}\beta}\right)+6x^{2}\left(\dfrac{x}{n\beta}\right)$
$=\frac{6x^{3}\left(\beta^{2}+5\beta+7\right)}{n\beta(\beta+1)(\beta+2)}+\frac{4x^{2}\beta+11}{n^{2}\beta(\beta+1)}+\frac{x}{n^{3}\beta}.$

### v2

@@ Abstract
This note aims to present a new sequence of positive linear operators involving the Wright function. Furthermore, the present research established the moments of these newly defined operators and estimated the convergence rate using the classical modulus of continuity. Additionally, the convergence rate in the Lipschitz spaces, their $A$-statistical convergence property have been covered.
@@
pacs:
MSC Classification] primary 41A25, 41A30, 41A36
@@ 1 Introduction
A new generalization of Bernstein operators was introduced periodically in approximation theory (see [1, 2, 3, 4, 5, 6]). Many positive linear operators are derived from special functions. For instance, the operators were introduced and studied in [7] using the Mittag-Leffler function, the operators were studied in [8] with the help of gamma function, and using Laguerre polynomials, the positive linear operators introduced by Gupta [9]. This note aims to merge two mathematical topics-approximation theory and special functions.
@@
named after the British mathematician E. M. Wright, has appeared for the first time in the case $\rho>0$ in connection with his investigations in the asymptotic theory of partitions in [10].
Note that $\phi(1,1;z)=I_{0}(2\sqrt{z})$,
where $I_{n}(z)=i^{-n}J_{n}(iz)=\dfrac{(z/2)^{n}}{\Gamma(1+n)}{}_{0}F_{1}\left(-;1+n;\dfrac{z^{2}}{4}\right)$, $n$ not a negative integer. The function $I_{n}(z)$ is called the modified Bessel function of the first kind of index $n$. Here, $J_{n}$ is Bessel function and $_{0}F_{1}$ is hypergeometric function. The modified function $I_{n}$ is related to $J_{n}$ in much the same way that the hyperbolic function is related to the trigonometric function.
@@
It was demonstrated in [11, 12] that, the modified Bessel function $I_{n}(z)$ of the first kind does not have zeros in the right half-plane. [13] contains an analysis of $n$-zeros of the function $I_{n}(z)$ in the left half-plane. Since $I_{n}(z)$ has no zeros in $0,\infty)$, the Wright function $\phi_{1,m}(z)$ also has no zeros in $0,\infty)$.
@@
$W_{n}^{(\beta)}(f;x)=\dfrac{1}{\phi_{1,\beta}\left(nx\right)}\sum_{k=0}^{\infty}f\left(\dfrac{k+\beta}{n}\right)\dfrac{(nx)^{k}}{k!\Gamma(k+\beta)},$ (1.2)
where $f\in E:=\left\{f\in C\left[0,\infty\right):\lim_{x\to\infty}\dfrac{f(x)}{1+x^{2}}\textrm{ is finite }\right\}$ and $C\left[0,\infty\right)$ denote the space of continuous functions defined on $0,\infty)$. Recall that the Banach lattice $E$ is endowed with the norm
@@
Research on approximating continuous signals using a sequence of positive linear operators is ongoing. We provide other findings along the same lines in this note, but with the aid of recently established positive linear operators $W_{n}^{(\beta)}$. There is no prior literature that introduced these operators (1.2). We estimate the moments and central moments of the operators (1.2) up to the fourth order in the second section. The rate of convergence of these positive linear operators is discussed in the next section. In the same section 3, we have proved that $W_{n}^{(\beta)}$ maps $E$ into itself. We have also determined the rate of convergence for these operators by utilizing the modulus of continuity. For the operators $W_{n}^{(\beta)}$, we have established a statistical Voronovskaya-type theorem in section 4.
@@ 2 Some Lemmas
Lemma 1.
[14, Theorem 6.1] Let $\alpha,\beta>0$. Then the following assertions are true:
@@
    $\left|W_{n}^{(\beta)}(t;x)-x\right|\leq\dfrac{\beta}{n};$
@@
    $\left|W_{n}^{(\beta)}(t^{2};x)-x^{2}\right|\leq\dfrac{x(1+2\beta)}{\beta n}+\dfrac{\beta^{2}}{n^{2}}.;$
@@
    $\left|W_{n}^{(\beta)}(t^{3};x)-x^{3}\right|\leq\dfrac{3x^{2}}{n\beta}+\dfrac{x(1+3\beta+\beta^{2})}{\beta n^{2}}+\dfrac{\beta^{3}}{n^{3}};$
@@
    $\left|W_{n}^{(\beta)}(t^{4};x)-x^{4}\right|\leq\frac{(4\beta+6)x^{3}}{n\beta(\beta+1)(\beta+2)}+\frac{(6\beta^{2}+12\beta+7)x^{2}}{n^{2}\beta(\beta+1)}+\frac{(4\beta^{3}+6\beta^{2}+4\beta+1)x}{n^{3}\beta}+\frac{\beta^{4}}{n^{4}}.$
Proof.
Since,
@@
$W_{n}^{(\beta)}(t;x)$ $=\dfrac{1}{\phi_{1,\beta}\left(nx\right)}\sum_{k=0}^{\infty}\dfrac{1}{n}\dfrac{(nx)^{k}}{(k-1)!\Gamma(k+\beta)}+\dfrac{\beta}{n}$
$=\dfrac{x}{\phi_{1,\beta}\left(nx\right)}\sum_{k=0}^{\infty}\dfrac{(nx)^{k}}{k!\Gamma(k+\beta+1)}+\dfrac{\beta}{n}=x\dfrac{\phi_{1,\beta+1}(nx)}{\phi_{1,\beta}(nx)}+\dfrac{\beta}{n}\leq x+\dfrac{\beta}{n}\textrm{ for }\beta>1.$
@@
$W_{n}^{(\beta)}(t^{2};x)$ $=\dfrac{1}{\phi_{1,\beta}\left(nx\right)}\sum_{k=1}^{\infty}\dfrac{(k+\beta)^{2}}{n^{2}}\dfrac{(nx)^{k}}{(k-1)!\Gamma(k+\beta)}$
$=\frac{1}{n^{2}\,\phi_{1,\beta}(nx)}\Bigg[(nx)^{2}\sum_{k=0}^{\infty}\frac{(nx)^{k}}{k!\,\Gamma(k+2+\beta)}+(1+2\beta)\,nx\sum_{m=0}^{\infty}\frac{(nx)^{m}}{m!\,\Gamma(m+1+\beta)}+\beta^{2}\,\phi_{1,\beta}(nx)\Bigg]$
$=\frac{1}{n^{2}\,\phi_{1,\beta}(nx)}\Big[(nx)^{2}\,\phi_{1,\beta+2}(nx)+(1+2\beta)\,nx\,\phi_{1,\beta+1}(nx)+\beta^{2}\,\phi_{1,\beta}(nx)\Big].$
@@
$\left|W_{n}^{(\beta)}(t^{2};x)-x^{2}\right|\leq\dfrac{x(1+2\beta)}{\beta n}+\dfrac{\beta^{2}}{n^{2}}.$
@@
$W_{n}^{(\beta)}(t^{3};x)$ $=$ $\frac{1}{n^{3}\phi_{1,\beta}(nx)}\sum_{k=0}^{\infty}(k+\beta)^{3}\frac{(nx)^{k}}{k!,\Gamma(k+\beta)}$
$=$ $\frac{1}{n^{3}\phi_{1,\beta}(nx)}\Big[(nx)^{3}\phi_{1,\beta+3}(nx)+3(1+\beta)(nx)^{2}\phi_{1,\beta+2}(nx)$
$+(1+3\beta+3\beta^{2})nx\phi_{1,\beta+1}(nx)+\beta^{3}\phi_{1,\beta}(nx)\Big].$
$=$ $x^{3}\dfrac{\phi_{1,\beta+3}\left(nx\right)}{\phi_{1,\beta}\left(nx\right)}+3(1+\beta)x^{2}\dfrac{\phi_{1,\beta+2}\left(nx\right)}{n\phi_{1,\beta}\left(nx\right)}+x(1+3\beta+\beta^{2})\dfrac{\phi_{1,\beta+1}\left(nx\right)}{n^{2}\phi_{1,\beta}\left(nx\right)}+\dfrac{\beta^{3}}{n^{3}}$
$\leq$ $\left|\dfrac{x^{3}\Gamma(\beta)}{\Gamma(\beta+3)}\right|+\left|\dfrac{3(1+\beta)x^{2}\Gamma(\beta)}{n\Gamma(\beta+2)}\right|+\left|\dfrac{x(1+3\beta+\beta^{2})\Gamma(\beta)}{n^{2}\Gamma(\beta+1)}\right|$
$=$ $\dfrac{x^{3}}{\beta(\beta+1)(\beta+2)}+\dfrac{3x^{2}}{n\beta}+\dfrac{x(1+3\beta+\beta^{2})}{\beta n^{2}}+\dfrac{\beta^{3}}{n^{3}}.$
@@
$\left|W_{n}^{(\beta)}(t^{3};x)-x^{3}\right|\leq\dfrac{3x^{2}}{n\beta}+\dfrac{x(1+3\beta+\beta^{2})}{\beta n^{2}}+\dfrac{\beta^{3}}{n^{3}}.$
@@
$W_{n}^{(\beta)}(t^{4};x)$ $=$ $\dfrac{1}{\phi_{1,\beta}\left(nx\right)}\sum_{k=0}^{\infty}\dfrac{(k+\beta)^{4}}{n^{4}}\dfrac{(nx)^{k}}{k!\Gamma(k+\beta)}$
$=$ $\frac{1}{n^{4}\phi_{1,\beta}(nx)}\Big[(nx)^{4}\phi_{1,\beta+4}(nx)+(4\beta+6)(nx)^{3}\phi_{1,\beta+3}(nx)$
$+(6\beta^{2}+12\beta+7)(nx)^{2}\phi_{1,\beta+2}(nx)+(4\beta^{3}+6\beta^{2}+4\beta+1)nx\phi_{1,\beta+1}(nx)+\beta^{4}\phi_{1,\beta}(nx)\Big].$
$=$ $x^{4}\dfrac{\phi_{1,\beta+4}\left(nx\right)}{\phi_{1,\beta}\left(nx\right)}+(4\beta+6)x^{3}\dfrac{\phi_{1,\beta+3}\left(nx\right)}{n\phi_{1,\beta}\left(nx\right)}+(6\beta^{2}+12\beta+7)x^{2}\dfrac{\phi_{1,\beta+2}\left(nx\right)}{n^{2}\phi_{1,\beta}\left(nx\right)}$
$+(4\beta^{3}+6\beta^{2}+4\beta+1)x\dfrac{\phi_{1,\beta+1}\left(nx\right)}{n^{3}\phi_{1,\beta}\left(nx\right)}+\dfrac{\beta^{4}}{n^{4}}$
$\leq$ $\left|\dfrac{x^{4}\Gamma(\beta)}{\Gamma(\beta+4)}\right|+\left|\dfrac{(4\beta+6)x^{3}\Gamma(\beta)}{n\Gamma(\beta+3)}\right|+\left|\dfrac{(6\beta^{2}+12\beta+7)x^{2}\Gamma(\beta)}{n^{2}\Gamma(\beta+2)}\right|$
$+\left|\dfrac{(4\beta^{3}+6\beta^{2}+4\beta+1)x\Gamma(\beta)}{n^{3}\Gamma(\beta+1)}\right|+\dfrac{\beta^{4}}{n^{4}}$
$=$ $\dfrac{x^{4}}{\beta(\beta+1)(\beta+2)(\beta+3)}+\dfrac{(4\beta+6)x^{3}}{n\beta(\beta+1)(\beta+2)}+\dfrac{(6\beta^{2}+12\beta+7)x^{2}}{n^{2}\beta(\beta+1)}$
$+\dfrac{(4\beta^{3}+6\beta^{2}+4\beta+1)x}{n^{3}\beta}+\dfrac{\beta^{4}}{n^{4}}.$
@@
$\left|W_{n}^{(\beta)}(t^{4};x)-x^{4}\right|\leq\dfrac{(4\beta+6)x^{3}}{n\beta(\beta+1)(\beta+2)}+\dfrac{(6\beta^{2}+12\beta+7)x^{2}}{n^{2}\beta(\beta+1)}+\dfrac{(4\beta^{3}+6\beta^{2}+4\beta+1)x}{n^{3}\beta}+\dfrac{\beta^{4}}{n^{4}}.$
Lemma 3.
@@
    $W_{n}^{(\beta)}(\phi_{x}^{2};x)\leq\frac{x(1+2\beta)}{\beta n}+\frac{2x\beta}{n}+\frac{\beta^{2}}{n^{2}}.$
@@
    $W_{n}^{(\beta)}(\phi_{x}^{3};x)\leq\frac{3x^{2}}{n\beta}+\frac{x(1+3\beta+\beta^{2})}{\beta n^{2}}+\frac{\beta^{3}}{n^{3}}+3x\left(\frac{x(1+2\beta)}{\beta n}+\frac{\beta^{2}}{n^{2}}\right)+\frac{3x^{2}\beta}{n}.$
@@
    $W_{n}^{(\beta)}(\phi_{x}^{4};x)\leq\left(\frac{4\beta+6}{\beta(\beta+1)(\beta+2)}+\frac{18}{\beta}+12+4\beta\right)\frac{x^{3}}{n}+\left(\frac{6\beta^{2}+12\beta+7}{\beta(\beta+1)}+\frac{4}{\beta}+12+4\beta+6\beta^{2}\right)\frac{x^{2}}{n^{2}}+\left(4\beta^{3}+4\beta^{2}+6\beta+4+\frac{1}{\beta}\right)\frac{x}{n^{3}}+\frac{\beta^{4}}{n^{4}}.$
Proof.
Using Lemma 2 and simple computation gives $W_{n}^{(\beta)}(\phi_{x}^{1};x)=0$ and
$W_{n}^{(\beta)}(\phi_{x}^{2};x)\leq\left|W_{n}^{(\beta)}(t^{2};x)-x^{2}\right|+2x\left|W_{n}^{(\beta)}(t;x)-x\right|+x^{2}\left|W_{n}^{(\beta)}(1;x)-1\right|\leq\frac{x(1+2\beta)}{\beta n}+\frac{2x\beta}{n}+\frac{\beta^{2}}{n^{2}}.$
@@
$\leq\left(\frac{4\beta+6}{\beta(\beta+1)(\beta+2)}+\frac{12}{\beta}+\frac{6(1+2\beta)}{\beta}+4\beta\right)\frac{x^{3}}{n}$
$+\left(\frac{6\beta^{2}+12\beta+7}{\beta(\beta+1)}+\frac{4(1+3\beta+\beta^{2})}{\beta}+6\beta^{2}\right)\frac{x^{2}}{n^{2}}+\left(\frac{4\beta^{3}+6\beta^{2}+4\beta+1}{\beta}+4\beta^{3}\right)\frac{x}{n^{3}}+\frac{\beta^{4}}{n^{4}}.$